OpenStudy (linn99123):

What is sin B? A.) 3/4 D.)4/3 C.)3/5 D.)4/5

OpenStudy (shiraz14):

Where's your diagram?

OpenStudy (shiraz14):

D - 4/5

OpenStudy (haseeb96):

\[\sin \theta = \frac{ perpendicualr }{ hypotenuse }\]

OpenStudy (haseeb96):

according to the figure the hypotenuse is 5 and perpendicular is 3 now put it in the sin function let theta = B

OpenStudy (haseeb96):

C . 3/5 is the correct answer